OpenStudy (anonymous):

Increase $200 by 10% and then decrease the new number by 10%. Is the final price same as the original one.

OpenStudy (lucaz):

increase by 10% is the same as multiply by 100%+10%=110%=1.1 200*1.1 then decrease by 10% is the same as multiply by 100%-10%=90%=0.9 so you have ($200*1.1)(0.9)=$198

OpenStudy (anonymous):

on Rs 100 the increased price would be Rs 110 on Rs 200 it would be 110/100*200 = Rs 220 now when 10 % is decreased from the new number on Rs 100 the decreased price would be Rs 90 on Rs 220 the decreased price would be Rs 90/100*220 = Rs 198 so the prices are not same